OpenStudy (anonymous):

A car dealer offers you two deals on a car that costs $16,000.Please calculate the monthly payment, given these two payment options the car dealer is offering. Payment Option 1: You can finance the car for 60 months with no interest if you make a $1,000 down payment. Payment Option 2: You can finance the car for 72 months (6 years) with 2% simple annual interest and no down payment. (Hint: To calculate simple annual interest, use the formula Interest = Principal * Rate * Time. Add the amount of interest to the price of the car.)

OpenStudy (anonymous):

Which monthly payment amount is lower? Please explain how you arrived at your answer and show all of your work. Where do I even start in this? It's so overwhelming...

OpenStudy (anonymous):

@bakonloverk

OpenStudy (bakonloverk):

its confusing me too

OpenStudy (anonymous):

LOL I broke @bakonloverk

OpenStudy (bakonloverk):

lol theres too many...words

OpenStudy (bakonloverk):

i know how to do it, i just dont know where to start

OpenStudy (anonymous):

^^^ It's alright lol. How many people does it take to answer this question we will never know... call reinforcements. lol

OpenStudy (bakonloverk):

lol @Notamathgenius

OpenStudy (notamathgenius):

@amistre64

OpenStudy (amistre64):

well, no interest and 1000 down what is: (16000 - 1000), divided by 60 ?

OpenStudy (amistre64):

2% simple interest a year amounts to 16000*.02/12 each month for 72 months

OpenStudy (anonymous):

16000-1000 is 15000. divided by 60 that is 250

OpenStudy (amistre64):

then for option 1, you are paying 250 a month

OpenStudy (anonymous):

and you got that from a formula correct?

OpenStudy (amistre64):

i believe so. no interest just means you are paying off the amount of the loan itself since you make a 1000 downpayment, you are only borrowing 15000 over the course of 60 months. 250 a month will payback the loan

OpenStudy (amistre64):

now, simple interest for 72 months is 6 years of interest 16000 * .02 * 6 is the amount if interest you would have to pay for option 2 (16000 + interest)/72 would be the payments for option 2

OpenStudy (anonymous):

for 16000 * .02 * 6 I got 19,200 Is that right?

OpenStudy (amistre64):

thats the total amount of interest for 6 years yes add that to the amount of the loan to get the total amount to be paid in 6 years

OpenStudy (anonymous):

16,000+19,200= 35,200. Then divide that by 72 is apparently, 488.888889 But I'm guessing I will need to round.

OpenStudy (amistre64):

248.888888 , but yes we could also look at this from: interest paid per month + principal paid per month\[16000(\frac{.02}{12})+16000\frac{1}{72}\] \[16000(\frac{.02}{12}+\frac{1}{72})\] \[16000(\frac{.12+1}{72})\] \[16000(\frac{1.12}{72})\]

OpenStudy (amistre64):

16000+16000*.02*6 = 17920 17920/72 = 248.888888 http://www.wolframalpha.com/input/?i=%2816000%2B16000*.02*6%29%2F72

OpenStudy (amistre64):

i see its 1920, NOT 19,200 your interest was over calculated the first time

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=16000*.02*6

OpenStudy (anonymous):

OH. See that]s probably why I got a totally different answer from what your showing me here, I understand what I did xD

OpenStudy (amistre64):

:) so option 1 = 250 a mnth option 2 = 249 a mnth

OpenStudy (anonymous):

So option 2 is lower :3 Thanks a lot! Really did help...

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

Thanks a lot :D